OpenStudy (anonymous):

Which of the following ordered pairs is not a solution to the inequality y greater than or equal to −4x − 2? Can someone help me?

OpenStudy (anonymous):

\[ y\geq -4x-2 \]

OpenStudy (anonymous):

for a point like \((\color{red}1,\color{blue}2)\), you would substitute:\[ \color{blue}2 \geq -4(\color{red}1)-2\\ 2\geq -4-2\\ 2\geq -6 \]

OpenStudy (anonymous):

Since the final statement is true, then \((1,2)\) is a solution.

OpenStudy (anonymous):

Do you get it?

OpenStudy (anonymous):

Kinda there are points that are given and that is not one like that so I am confused

OpenStudy (anonymous):

The closed one is (1,3)

OpenStudy (anonymous):

Okay, can you try to do what I did, but with the points they gave?

OpenStudy (anonymous):

yeah that makes sense